OpenStudy (anonymous):

how do you find point of intersection y=x^2 and y=1/x without using the graph ?

OpenStudy (turingtest):

\[x^2=\frac1x\]etc. solve for x

OpenStudy (anonymous):

\[x^2=y=\frac{1}{x}\]

OpenStudy (anonymous):

\[x^3=1\] \[x=1\]

OpenStudy (anonymous):

thanks.How would you find it for y=x^2 , y=x^3

OpenStudy (anonymous):

\[x^2=x^3\] divide both sides by \[x^2\]

OpenStudy (anonymous):

@yosucka ...please do not completely solve problems for someone. TuringTest's hint was a great starting point, but you made it so no effort at all was required on the part of the asker.